A can do a piece of work in 12 days and B can do a piece of work in 18 days. They started working together but A left 3 days before the completion of the work. The total number of days to complete the work is ??
Answer by dkppathak(439)

A can do a piece of work in 12 days
A one day work is =1/12
B can do a piece of work in 18 days
B one day work is =1/18
A and B both together work for one day= 1/12+1/18
=5/36 part
let total work completed in x days
both work for (x-3) and B work alone 3 days
both work for (x-3) days =(x-3)5/36=5x-15/36
B alone three days work =3x1/18=1/6
as per give conditions
1/6+ 5x-15/36 =1
by LCM solving we will get
-9+5x/36=1
-9+5x=36
5x =45
x=45/5=9 days
answer work completed in 9 days